In how many arrangments can 5 women and 5 men be placed in a line if all men line up first?
-please show work

Answer by jim_thompson5910(35256)

5! = 5*4*3*2*1 = 120 ways to arrange the 5 men.

5! = 120 ways to arrange the 5 women.

So 120*120 = 14400 different ways to arrange the 10 people.